# Discontinuous integro-differential equations and sliding mode control

### v2

@@ 4.1 SMC for a distributed input delay system
To compare the SMC for IDE with the SMC for ODE , we also present the simulation results of the system (45)-(47) with $\Phi=\mathbf{0}$ (please see Figures 5, 6 and 7). The simulations show a faster convergence of the delay-free ($\Phi=\mathbf{0}$) system to zero.
@@
4.2 SMC for dynamical systems in Banach spaces
Let us consider a control system modeled by a differential equation in a Banach space [DaleckyKrein1974:Book], [Yosida1980:Book], [Pazy1983:Book]
$y(t)=Cx(t)\in\mathbb{R}^{k},\vskip-2.84526pt$ (64)
where $C:\mathbb{B}\to\mathbb{R}^{k}$ is a linear bounded operator. For any $u\in L^{1}_{\rm loc}([t^{0},+\infty);\mathbb{R}^{m})$, the open-loop systems (63) is well-posed and has the so-called mild solution [Pazy1983:Book, page 106]
where the integral is understood in the sense of Bochner [Yosida1980:Book]. The mild solution $x:[t^{0},t^{0}+T)\to\mathbb{B}$ with $0<T\leq+\infty$ is the classical (resp., strong [Pazy1983:Book, page 109]) solution of (63) if $x(t)\!\in\!\mathcal{D}(A)$ satisfies (63) for (almost) all $t\in[t^{0},t^{0}+T)$.
@@ 4.2 SMC for dynamical systems in Banach spaces
where $\tilde{u}$ is a measurable function. Even if a function $\tilde{u}$ is continuous in $y$, the non-linearity in the right-hand side of the closed-loop system may not be locally Lipschitz continuous, so, the conventional theory of nonlinear abstract differential equations (see, e.g., [Pazy1983:Book], [EngelNagel2000:Book]) is not applicable for its analysis. In the case of SMC design [Orlov_elal2004:IJC], the closed-loop system (63), (64), (66) is ill-posed. Let us regularize the system (63), (64) following the concept of Utkin solutions [Orlov2008:Book], [PolyakovFridman2014:JFI].
@@
In other words, an Utkin solution $x$ of (63), (64), (66) is a mild solution with a properly selected $u(t)\!\in\!U(t,Cx(t))$. If $B$ maps $\mathbb{R}^{m}$ into the domain of $A$ (i.e., $\mathrm{range}(B)\!\subset\!\mathcal{D}(A)$), then, by [Pazy1983:Book, Theorem 2.9, Chapter 4], any mild solution is a strong solution. By [Pazy1983:Book, Theorem 2.4, Chapter 1], the function
@@
is continuous. Moreover, taking into account $\frac{d\Psi(s)B}{ds}=\Psi(s)AB$ and $\frac{d\Psi(s)x^{0}}{ds}=\Psi(s)Ax^{0},\forall s>0,\forall x^{0}\in\mathcal{D}(A)$ (see, [Pazy1983:Book, Theorem 2.4, Chapter 1]), we derive
$\begin{split}\tfrac{d}{dt}y(t)=\tfrac{d}{dt}Cx(t)=&\frac{d}{dt}C\Psi(t-t)x^{0}\\ &+\frac{d}{dt}\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)B}_{\in\mathbb{R}^{k\times m}}(u(\tau)+\gamma(\tau))d\tau\\ =&C\Psi(t-t)Ax^{0}+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)AB}_{\Phi(t-\tau)}(u(\tau)+\gamma(\tau))d\tau\end{split}$
@@
$\dot{y}(t)=p(t-t^{0})+CB(u(t)+\gamma(t))+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (67)
@@
Let $y_{\rm eq}$ be an Utkin solution of the IDE (67), (68) with the feedback law (66), and let $u_{\rm eq}$ be the corresponding equivalent control. By Theorem 19, the pair $(y,u_{\rm eq})$ always exists if $\tilde{u}$ is a locally essentially bounded measurable function. If $x_{\rm eq}$ is a mild solution of (63) given by (65) with $u\!=\!u^{\rm eq}$, then
$\begin{split}\tfrac{d}{dt}Cx_{\rm eq}(t)=&C\Psi(t-t)Ax^{0}+CB(u_{\rm eq}(t)+\gamma(t))\\ &+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u_{\rm eq}(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$
@@
$\begin{split}-\xi(t,t^{0},x^{0})=&\,\Phi(t-t^{0})\xi(t^{0},t^{0},x^{0})-\Phi(0)\xi(t,t^{0},x^{0})\\ =&\int^{t-t^{0}}_{0}\frac{d}{d\sigma}\left(\Phi(\sigma)\xi(t-\sigma,t^{0},x^{0})\right)d\sigma\\ =&\int^{t-t^{0}}_{0}C\Psi(\sigma)A^{2}B\xi(t-\sigma,t^{0},x^{0})d\sigma\\ &-\int^{t-t^{0}}_{0}\Phi(\sigma)\dot{\xi}(t-\sigma,t^{0},x^{0})d\sigma\end{split}\vskip-2.84526pt$
@@
$\begin{split}\dot{y}(t,t^{0},x^{0})\in&\,p(t,t^{0},x^{0})+CB(U(t,y(t,t^{0},x^{0}))+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t-\tau)\gamma(\tau)d\tau+\xi(t,t^{0},x^{0})\\ &+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau,\\ \dot{\xi}(t,t^{0},x^{0})\in&\,U(t,y(t,t^{0},x^{0})),\quad\xi(t^{0})=0,\quad y(t^{0})=Cx^{0},\end{split}\vskip-2.84526pt$
@@
1) Regularization of PDE using IDE. The example (1)–(6) admits the representation (63) with $A\!=\!\nu\frac{\partial^{2}}{\partial z^{2}}$, $D(A)\!=\!H^{2}((0,1);\mathbb{R})\cap H^{1}_{0}((0,1);\mathbb{R})$, $(Bx)(z)\!=\!\beta(z)x(z)$, where $H^{2}$, $H^{1}_{0}$ are Sobolev spaces. The closed densely defined operator $A:\mathcal{D}(A)\subset L^{2}((0,1);\mathbb{R})\to L^{2}((0,1);\mathbb{R})$ is a generator of the strongly continuous semigroup of linear bounded operators on $L^{2}((0,1);\mathbb{R})$ (see, e.g., [Pazy1983:Book, page 212]). The output (64) is given by $y(t)=Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$. The control law (66) is defined by the formula (6). Since $\beta\in\mathcal{D}(A)$, then, by Corollary 6, the system (1)–(6) has an Utkin solution $x(t)\!\in\!\mathcal{D}(A)$, $\forall x^{0}\in\mathcal{D}(A)$, $\forall q\!\in\!L^{\infty}((t^{0},+\infty);\mathbb{R})$ in (6).
2) SMC design using PDE-based approach. The closed-loop heat control system (1)-(6) is forward complete, i.e., all Utkin solutions exist globally in time. Indeed, considering $V(t)\!=\!\int^{1}_{0}\!x^{2}(t,z)dz$ and using integration by parts, we derive
$\dot{V}(t)=2\int^{1}_{0}\nu x(t,z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}+x(t,z)\beta(z)u_{\rm eq}(t)dz\vskip-5.69054pt$
$2\int^{1}_{0}\nu\left(\tfrac{\partial x(t,z)}{\partial z}\right)^{2}+x(t,z)\beta(z)u_{\rm eq}(t)dz,\vskip-2.84526pt$
@@
$\dot{V}(t)\leq-\nu 2\pi^{2}V(t)+2\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}\sqrt{V(t)},$
@@
$\limsup_{t\to+\infty}\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\|q\|_{L^{\infty}}\|\beta\|_{L^{2}}}{\nu\pi^{2}}.\vskip-2.84526pt$ (69)
@@
Taking into account that $u_{\rm eq}(t)\in-q(t)\overline{\mathrm{sign}}(Cx(t))$, we conclude that, for $Cx(t)\!\neq\!0$, it holds $u_{\rm eq}(t)\!=\!-q(t)\mathrm{sign}(Cx(t))$,
@@
almost everywhere. If $CB=\int^{1}_{0}\xi(z)\beta(z)dz\neq 0$ then, taking $\delta>0$ and $q(t)=-\frac{\rho}{CB}$ with $\rho>0$, we have due to (69)):
$\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{\rho\|\beta\|_{L^{2}}}{\pi^{2}|CB|}+\delta\vskip-2.84526pt$
@@
$\rho>\tfrac{\pi^{2}|CB|\left(\bar{\gamma}|CB|+\nu\delta\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\right)}{\pi^{2}\left|CB\right|-\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}},\quad\bar{\gamma}\geq\|\gamma\|_{L^{\infty}},\vskip-2.84526pt$
@@
$\tfrac{d|Cx(t)|}{dt}\!\leq\!\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}\!+\!\lambda\xi\right\|_{L^{2}}\!\left(\tfrac{\rho\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}\!+\!\delta\right)\!+|CB|\cdot\|\gamma\|_{\infty}-\rho\!<\!0\;$
@@
3) SMC design using IDE-based approach. The same result can be obtained considering the input-output dynamics of the heat control system. Using the modal decomposition
@@
where $\phi_{i}(z)=\sqrt{2}\sin(\sqrt{\lambda_{i}}z),\lambda_{i}=\pi^{2}i^{2}$, the system (1) - (3) can be rewritten (see, e.g. [Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC]) in the form
@@
$p(t,t^{0})=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-t^{0})}x_{i}^{0}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz,\vskip-5.69054pt$
$\Phi(t,\tau)=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-\tau)}b_{i}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz.\vskip-2.84526pt$
Since $|\Phi(t,\tau)|\leq\nu e^{-\nu\pi^{2}(t-\tau)}\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}$ then the inequality (71) yields $M=\frac{\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|\Phi(t,\tau)|d\tau}{|CB|}\!<\!1$. Let $q(t)=-\frac{|CB|\cdot\|\gamma\|_{L_{\infty}}+|p(t,t^{0})|+M\|\gamma\|_{L^{\infty}}+\delta}{(1-M)CB}$ with $\delta>0$. By Theorem 20 and Corollary 4, the system (76), (6) has the unique Utkin solution and $y(t)=0$ for all $t\geq T$, where $T\geq t^{0}$ is some instant of time. This means that the Utkin solution of the system (1)–(6) satisfies (73) for all $t\geq T$.
4) Stability of the heat system in the sliding mode. Let us show that the system (72), (73), (3) is globally exponentially stable. Considering again the functional $t\to V(t)=\int^{1}_{0}x^{2}(t,z)dz$ , we derive
@@
The identity (73) gives
$\begin{split}\left|\int^{1}_{0}\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\right|=&\left|\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\right|\\ \leq&\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\sqrt{V(t)}.\vskip-2.84526pt\end{split}$
@@
$\dot{V}(t)\leq 2V(t)\left(\tfrac{\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\|\beta\|_{L^{2}}}{|CB|}-\pi^{2}\right)<0\text{ for }V(t)\neq 0,\vskip-2.84526ptx$
i.e., the system (72), (73), (3) is globally exponentially stable.
@@
In this case, we have $CB\approx 0.33$, $\|\beta\|_{L^{2}}\approx 0.55$ and $\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\approx 0.8$ for $\lambda=\pi^{2}$. The figures 8, 9, 10 show simulation results for the heat system (1)-(5) with the sliding mode control (6), $q(t)=-\frac{\rho}{CB},\rho=1$ and the matched perturbation $\gamma(t)=0.5$. The initial state $x^{0}\in\mathcal{D}(A)$ is defined as follows $x_{0}(z)=10z(1-z),z\in[0,1]$. The method of modal decomposition [Balas1988:JMAA], [PrieurTrelat2019:TAC], [KatzFridman2021:EJC], [Ayamou_etal2025:IMA] (with 60 modes) is utilized for numerical simulation of the closed-loop SMC system. The time sampling is $0.001$.
@@
The finite-dimensional IDE-based procedure of regularization of infinite-dimensional systems with sliding mode control can be extended to other systems using, for example, the methodology suggested in [Polyakov2025:CDC]. Further study of the SMC design for such systems goes out of the scope of this paper.

### v3

@@ 4.1 SMC for a distributed input delay system
4.2 SMC for dynamical systems in Banach spaces
Let us consider a control system modeled by a differential equation in a Banach space [12], [51], [40]
$y(t)=Cx(t)\in\mathbb{R}^{k},\vskip-2.84526pt$ (65)
where $C:\mathbb{B}\to\mathbb{R}^{k}$ is a linear bounded operator. For any $u\in L^{1}_{\rm loc}([t^{0},+\infty);\mathbb{R}^{m})$, the open-loop systems (64) is well-posed and has the so-called mild solution [40, page 106]
where the integral is understood in the sense of Bochner [51]. The mild solution $x:[t^{0},t^{0}+T)\to\mathbb{B}$ with $0<T\leq+\infty$ is the classical (resp., strong [40, page 109]) solution of (64) if $x(t)\!\in\!\mathcal{D}(A)$ satisfies (64) for (almost) all $t\in[t^{0},t^{0}+T)$.
@@ 4.2 SMC for dynamical systems in Banach spaces
where $\tilde{u}$ is a measurable function. Even if a function $\tilde{u}$ is continuous in $y$, the non-linearity in the right-hand side of the closed-loop system may not be locally Lipschitz continuous, so, the conventional theory of nonlinear abstract differential equations (see, e.g., [40], [15]) is not applicable for its analysis. In the case of SMC design [35], the closed-loop system (64), (65), (67) is ill-posed. Let us regularize the system (64), (65) following the concept of Utkin solutions [37], [43].
@@
In other words, an Utkin solution $x$ of (64), (65), (67) is a mild solution with a properly selected $u(t)\!\in\!U(t,Cx(t))$.
If $\mathrm{range}(B)\!\subset\!\mathcal{D}(A)$, then, by [40, Theorem 2.9, Chapter 4], any mild solution is a strong solution. By [40, Theorem 2.4, Chapter 1], the function
@@
is continuous. Moreover, taking into account $\frac{d\Psi(s)B}{ds}=\Psi(s)AB$ and $\frac{d\Psi(s)x^{0}}{ds}=\Psi(s)Ax^{0},\forall s>0,\forall x^{0}\in\mathcal{D}(A)$ (see, [40, Theorem 2.4, Chapter 1]), we derive
$\small\begin{split}\tfrac{d}{dt}y(t)=\tfrac{d}{dt}Cx(t)=&\frac{d}{dt}C\Psi(t-t)x^{0}\\ &+\frac{d}{dt}\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)B}_{\in\mathbb{R}^{k\times m}}(u(\tau)+\gamma(\tau))d\tau\\ =&C\Psi(t-t)Ax^{0}+CB(u(t)+\gamma(t))\\ &+\int^{t}_{t^{0}}\underbrace{C\Psi(t-\tau)AB}_{\Phi(t-\tau)}(u(\tau)+\gamma(\tau))d\tau\end{split}$
@@
$\dot{y}(t)=p(t-t^{0})+CB(u(t)+\gamma(t))+\int^{t}_{\sigma}\!\!\Phi(t,\tau)(u(\tau)+\gamma(\tau))d\tau,\vskip-2.84526pt$ (68)
@@
Let $y_{\rm eq}$ be an Utkin solution of the IDE (68), (69) with the feedback law (67), and let $u_{\rm eq}$ be the corresponding equivalent control. By Theorem 19, the pair $(y,u_{\rm eq})$ always exists if $\tilde{u}$ is a locally essentially bounded measurable function. If $x_{\rm eq}$ is a mild solution of (64) given by (66) with $u\!=\!u_{\rm eq}$, then
$\small\begin{split}\tfrac{d}{dt}Cx_{\rm eq}(t)=&C\Psi(t-t)Ax^{0}+CB(u_{\rm eq}(t)+\gamma(t))\\ &+\int^{t}_{\sigma}\!\!\Phi(t-\tau)(u_{\rm eq}(\tau)+\gamma(\tau))d\tau,\end{split}\vskip-2.84526pt$
@@
$\small\begin{split}-\xi(t,t^{0},x^{0})=&\,\Phi(t-t^{0})\xi(t^{0},t^{0},x^{0})-\Phi(0)\xi(t,t^{0},x^{0})\\ =&\int^{t-t^{0}}_{0}\frac{d}{d\sigma}\left(\Phi(\sigma)\xi(t-\sigma,t^{0},x^{0})\right)d\sigma\\ =&\int^{t-t^{0}}_{0}C\Psi(\sigma)A^{2}B\xi(t-\sigma,t^{0},x^{0})d\sigma\\ &-\int^{t-t^{0}}_{0}\Phi(\sigma)\dot{\xi}(t-\sigma,t^{0},x^{0})d\sigma\end{split}\vskip-2.84526pt$
@@
$\small\begin{split}\dot{y}(t,t^{0},x^{0})\in&\,p(t,t^{0},x^{0})+CB(U(t,y(t,t^{0},x^{0}))+\gamma(t))\\ &+\int^{t}_{t^{0}}\Phi(t-\tau)\gamma(\tau)d\tau+\xi(t,t^{0},x^{0})\\ &+\int^{t}_{t^{0}}C\Psi(t-\tau)A^{2}B\xi(\tau,t^{0},x^{0})d\tau,\\ \dot{\xi}(t,t^{0},x^{0})\in&\,U(t,y(t,t^{0},x^{0})),\qquad\xi(t^{0})=0,\qquad y(t^{0})=Cx^{0},\end{split}\vskip-2.84526pt$
@@
1) Regularization of PDE using IDE. The example (1)–(6) admits the representation (64) with $A\!=\!\nu\frac{\partial^{2}}{\partial z^{2}}$, $D(A)\!=\!H^{2}((0,1);\mathbb{R})\cap H^{1}_{0}((0,1);\mathbb{R})$, $(Bx)(z)\!=\!\beta(z)x(z)$, where $H^{2}$, $H^{1}_{0}$ are Sobolev spaces. The closed densely defined operator $A:\mathcal{D}(A)\subset L^{2}((0,1);\mathbb{R})\to L^{2}((0,1);\mathbb{R})$ is a generator of the strongly continuous semigroup of linear bounded operators on $L^{2}((0,1);\mathbb{R})$ (see, e.g., [40, page 212]). The output (65) is given by $y(t)=Cx(t)=\int^{1}_{0}\xi(z)x(t,z)dz$. The control law (67) is defined by the formula (6). Since $\beta\in\mathcal{D}(A)$, then, by Corollary 6, the system (1)–(6) has an Utkin solution $x(t)\!\in\!\mathcal{D}(A)$, $\forall x^{0}\in\mathcal{D}(A)$, $\forall q\!\in\!L^{\infty}((t^{0},+\infty);\mathbb{R})$ in (6).
2) SMC design using PDE. The closed-loop heat control system (1)-(6) is forward complete, i.e., all Utkin solutions exist globally in time. Indeed, considering $V(t)\!=\!\int^{1}_{0}\!x^{2}(t,z)dz$ and using integration by parts, we derive
$\vskip-2.84526pt\small\begin{split}\dot{V}(t)=&\,2\int^{1}_{0}\nu x(t,z)\tfrac{\partial^{2}x(t,z)}{\partial z^{2}}+x(t,z)\beta(z)(u_{\rm eq}(t)+\gamma(t))dz\\[-5.69054pt] =&-2\int^{1}_{0}\nu\left(\tfrac{\partial x(t,z)}{\partial z}\right)^{2}+x(t,z)\beta(z)(u_{\rm eq}(t)+\gamma(t))dz,\end{split}$
@@
$\dot{V}(t)\leq-\nu 2\pi^{2}V(t)+2(\|q\|_{L^{\infty}}+\|\gamma\|_{L^{\infty}})\|\beta\|_{L^{2}}\sqrt{V(t)},$
@@
$\limsup_{t\to+\infty}\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{(\|q\|_{L^{\infty}}+\|\gamma\|_{L^{\infty}})\|\beta\|_{L^{2}}}{\nu\pi^{2}}.\vskip-2.84526pt$ (70)
@@
Taking into account that $u_{\rm eq}(t)\in q(t)\overline{\mathrm{sign}}(Cx(t))$, we conclude that, for $Cx(t)\!\neq\!0$, it holds $u_{\rm eq}(t)\!=\!q(t)\mathrm{sign}(Cx(t))$,
@@
almost everywhere. If $CB=\int^{1}_{0}\xi(z)\beta(z)dz\neq 0$ then, taking $\delta>0$ and $q(t)=-\frac{\rho}{CB}$ with $\rho>0$, we have due to (70):
$\|x(t,\cdot)\|_{L^{2}}\leq\tfrac{(\rho+\|\gamma\|_{L^{\infty}}|CB|)\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}+\delta\vskip-2.84526pt$
@@
$\rho\!>\!\tfrac{|CB|\left(\pi^{2}\bar{\gamma}|CB|+\left(\nu\delta\pi^{2}+\bar{\gamma}\|\beta\|_{L^{2}}\right)\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\right)}{\pi^{2}\left|CB\right|-\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}},\quad\bar{\gamma}\!\geq\!\|\gamma\|_{L^{\infty}},\vskip-2.84526pt$
@@
$\small\vskip-2.84526pt\begin{split}\tfrac{d|Cx(t)|}{dt}\!\leq&\nu\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}\!+\!\lambda\xi\right\|_{L^{2}}\!\left(\tfrac{(\rho+\|\gamma\|_{L^{\infty}}|CB|)\|\beta\|_{L^{2}}}{\nu\pi^{2}|CB|}\!+\!\delta\right)\\ &+|CB|\cdot\|\gamma\|_{\infty}-\rho<0\end{split}$
@@
3) SMC design using IDE. The same result can be obtained considering the input-output dynamics of the heat control system. Using the modal decomposition
@@
where $\phi_{i}(z)=\sqrt{2}\sin(\sqrt{\lambda_{i}}z),\lambda_{i}=\pi^{2}i^{2}$, the system (1) - (3) can be rewritten (see, e.g. [5], [45], [22]) in the form
@@
$p(t{\color[rgb]{0,0,1}-}t^{0})=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-t^{0})}x_{i}^{0}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz,\vskip-5.69054pt$
$\Phi(t{\color[rgb]{0,0,1}-}\tau)=\nu\sum_{i=1}^{\infty}e^{-\nu\lambda_{i}(t-\tau)}b_{i}\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)\phi_{i}(z)dz.\vskip-2.84526pt$
Since $|\Phi(t{\color[rgb]{0,0,1}-}\tau)|\leq\nu e^{-\nu\pi^{2}(t-\tau)}\|\beta\|_{L^{2}}\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}$ then the inequality (72) yields $M=\frac{\operatorname*{ess\,sup}_{t\geq t^{0}}\int^{t}_{t^{0}}\|\Phi(t{\color[rgb]{0,0,1}-}\tau)|d\tau}{|CB|}\!<\!1$.
Let $q(t)=-\frac{|CB|\bar{\gamma}+\bar{p}+M\bar{\gamma}+\delta}{(1-M)CB}$ with $\delta>0$ and $\bar{p}\geq|p(t-t^{0})|$. By Theorem 20 and Corollary 4, the system (77), (6) has the unique Utkin solution and $y(t)=0$ for all $t\geq T$, where $T\geq t^{0}$ is some instant of time. This means that the Utkin solution of the system (1)–(6) satisfies (74) for all $t\geq T$.
4) Stability analysis of the heat system in the sliding mode. Let us show that the system (73), (74), (3) is globally exponentially stable. Considering again the functional $t\to V(t)=\int^{1}_{0}x^{2}(t,z)dz$ , we derive
@@
The identity (74) and Cauchy-Schwartz inequality give
$\small\vskip-2.84526pt\begin{split}\left|\int^{1}_{0}\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}x(t,z)dz\right|=&\left|\int^{1}_{0}\left(\tfrac{\partial^{2}\xi(z)}{\partial z^{2}}+\lambda\xi(z)\right)x(t,z)dz\right|\\ \leq&\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\sqrt{V(t)}.\vskip-2.84526pt\end{split}$
@@
$\dot{V}(t)\leq 2V(t)\left(\tfrac{\left\|\frac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\|\beta\|_{L^{2}}}{|CB|}-\pi^{2}\right)<0\,\text{ for }\,V(t)\neq 0.\vskip-2.84526pt$
So, the system (73), (74), (3) is globally exponentially stable.
@@
In this case, we have $CB\approx 0.33$, $\|\beta\|_{L^{2}}\approx 0.55$ and $\left\|\tfrac{\partial^{2}\xi}{\partial z^{2}}+\lambda\xi\right\|_{L^{2}}\approx 0.8$ for $\lambda=\pi^{2}$. The figures 8, 9, 10 show simulation results for the heat system (1)-(5) with the sliding mode control (6), $q(t)=-\frac{\rho}{CB},\rho=1$ and the matched perturbation $\gamma(t)=0.5$. The initial state $x^{0}\in\mathcal{D}(A)$ is defined as follows $x_{0}(z)=10z(1-z),z\in[0,1]$. The method of modal decomposition [5], [45], [22], [4] with 60 modes is utilized for numerical simulation of the closed-loop SMC system. The time sampling is $0.001$.
@@
The finite-dimensional IDE-based procedure of regularization of infinite-dimensional systems with sliding mode control can be extended to other systems using, for example, the methodology suggested in [44]. Being interesting future research topic, a further study of the SMC design for such systems goes out of the scope of this paper.